OpenStudy (anonymous):

A square pyramid has a total surface area of eight hundred square feet. The area of the base is four hundred square feet. What is the area of one triangular face of the lateral surface?

OpenStudy (anonymous):

Plz help me

OpenStudy (anonymous):

Area of the four lateral trianngular surfaces = 800-400 = 400sft. area of one lateral triangular surface = 400/100 =100sft

OpenStudy (anonymous):

are u sure that it wouldnt b 4

OpenStudy (anonymous):

u dont have to give the answer just how to do the question if u want

OpenStudy (phi):

Sparrow gave you the answer. Here's what he did. A square pyramid has 4 (slanted sides) plus a bottom. The surface area of the pyramid is the sum of the areas of the 4 sides and the bottom. He subtracted the area of the bottom from the total. total - bottom = 800 - 400 = 400. That leaves 400 square feet for the 4 sides. All 4 sides are identical because the pyramid is square. So each must have 100 sq ft. , because 100 times 4 = 400.

OpenStudy (anonymous):

thank you